The USD appreciates 4% versus the JPY. The MXN depreciates 4% versus the USD. What is the approximate appreciation or depreciation we might see in the JPY/MXN exchange rate?
Answer #1

Lets say there was initially 1USD versus JPY
After appreciation, the rate becomes 1.04 JPY/USD

Lets say there was initially 1USD versus MXN
After depreciation, the rate becomes 0.96 USD/MXN

Initial rate will be 1JPY/MXN
Afterwards, rate will be =1.04*0.96=0.9984 JPY/MXN

Hence, MXN depreciated by 1-0.9984=0.16%